# Inertial Coordination Games

### v2

@@ 1   Introduction
Our main result (Theorem 1) develops a tight connection between the speed of learning and limit play: the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically. Sub-quadratic learning includes canonical learning environments such as (i) independent and identically distributed signals; and (ii) the receipt of a one-time signal. Super-quadratic learning includes social learning where in each period agents are randomly matched and sample their partner’s information sets. The tight connection we develop between learning speeds and limit play has simple but sharp economic implications. Slow (sub-quadratic) learning implies history independence: the initial shock to aggregate play as well the fine details of what agents learn in the interim is irrelevant for limit play. Conversely, when learning speeds are fast (super-quadratic), this induces history dependence such that initial shocks can be self-fulfilling: whether shocks propagate or not is jointly sensitive to both the fundamentals and the magnitude of the shock. [^2] See Section I of Krugman (1991) for a discussion of history-dependence and expectations in economics.
@@
Our model of inertial coordination games imposes a lag between players’ actions and their effect on the time-varying state. Hence, when a player contemplates whether or not to play the risky action, his belief about the current state depends on his belief about past play. But that, in turn, depends on his belief about other players’ past beliefs which themselves are about play yet further back into the past. Thus, the path of past play (and beliefs)—even of the distant past—multiply forward to shape incentives in the present. [^3] Note that our environment is, in a loose sense, the opposite of coordination games with switching frictions (Frankel and Pauzner, 2000). In those models, switching frictions lead players to form beliefs about future play and hence future beliefs; by contrast, in our model, inertia leads players to form beliefs about past play and hence past beliefs. An important by-product of inertia is that it offers a simple yet novel means of obtaining unique predictions in coordination environments. At the same time, our model is motivated by—and consistent with—recent work documenting empirical facts about bank runs. Correia, Luck, and Verner (2024) find “most bank failures are the result of a deterioration in bank solvency” (in our model, the shock to initial play) rather than being purely self-fulfilling (predicted by static coordination games). Furthermore, they find that “depositors tend to be slow to react to an increased risk of bank failure” (in our model, the inertia). Hence, one of our contributions is to offer a simple dynamic model which reconciles empirical regularities about ‘real world’ coordination environments with the more abstract logic of learning and higher-order beliefs in coordination games. Further, our characterization of the connection between speed and limit play is invariant to the degree of inertia (duration of lag) and, as inertia vanishes, the equilibrium converges to an equilibrium of the model with purely contemporaneous incentives. [^4] This is made precise in Appendix B.
@@
In what sense does our model relate to the extant literature on coordination games? It turns out that static global games (Carlsson and Van Damme, 1993) are recovered exactly as a special case of our model in which players only receive a single signal at time $t=1$: since this learning process is sub-quadratic (posterior precisions are unchanged after $t=1$), the risk-dominant profile is—as in global games—played in the limit. This connection arises from observing that when there is no interim learning, best-responding to expected past play along each time step is equivalent in our model to the interim elimination of strictly dominated strategies in the static game. As we make precise in Online Appendix V, this uncovers a formal connection between “contagion” arguments over type spaces (which encode hierarchies of higher-order beliefs) and “contagion” over time (where the current state depends on past play). Thus, our results generalize the predictions of static global games to sub-quadratic learning and offer a simple dynamic learning foundation for the ‘vanishing noise limit’ commonly invoked in that literature.
@@
Finally, our model makes sharp qualitative predictions about the path of play (Proposition 3) whenever learning is slow enough such that the risk-dominant profile is played in the limit. In particular, transition dynamics toward limit play depend starkly on the precision of signals. When signals are precise, aggregate play exhibits a sudden phase transition from approximately all players playing the non-risk-dominant action (potentially for a long while) to approximately all players playing the risk-dominant action thereafter. Conversely, when signals are noisy, aggregate play exhibits a gradual transition as the measure of players playing risk-dominant action increases gradually. These differing transition regimes are driven by differences in time-varying heterogeneity in beliefs. The importance of heterogeneity has been recognized as a crucial determinant of equilibrium uniqueness in coordination games (Morris and Shin, 2006). Our results complement this insight by highlighting that the speed of learning—which drives the how the cross-sectional distribution of beliefs vary over time—is also a crucial determinant of the path of play. It also suggests that “spikes” in the time-series of aggregate actions (e.g., withdrawal decisions or adoption of networked goods) can be consistent with transition to limit equilibrium play and need not necessarily correspond to “equilibrium shifts” (Chwe, 2013; Morris and Yildiz, 2019).
@@
Related literature. Our paper directly relates to the large literature on global games. A broad takeaway from the literature on static global games (Carlsson and Van Damme, 1993; Kajii and Morris, 1997; Frankel, Morris, and Pauzner, 2003, among others) is that perturbing common knowledge by introducing a vanishing amount of private uncertainty often selects the risk dominant equilibrium. This approach has been widely deployed in a variety of economic applications to select among multiple equilibria (see, e.g., Morris and Shin (1998); Goldstein and Pauzner (2005) among many others). Our model nests the canonical global games model of Carlsson and Van Damme (1993) as a special case with a private signal at $t=1$ and no learning thereafter. The “contagion argument” in global games where players iteratively eliminate interim dominated strategies is sometimes viewed metaphorically or as a convenient way to select among equilibria. [^5] See, Morris (2002) and Chapter 2 of Spiegler (2024). Our simple dynamic environment with an endogeneously changing state thus offers an explicit dynamic foundation for the “contagion” logic of global games consistent with recently documented facts on bank runs (Correia, Luck, and Verner, 2024). Moreover, we substantially generalize the prediction of risk-dominant selection to the regime of sub-quadratic learning.
@@ 2   Model
Payoffs. There is a unit measure of agents indexed by $i\in[0,1]$. [^12] Endow $[0,1]$ with the Lebesgue extension constructed in Sun and Zhang (2009) so that the continuum law of large number holds. Time is discrete and indexed by $t\in\mathcal{T}=\{1,2,\ldots\}$. At each time $t\in\mathcal{T}$, each agent $i$ chooses between a risky action $(a_{it}=1)$ or a safe action $(a_{it}=0)$. Let $\lambda_{t}$ denote the measure of agents who play the risky action at time $t$.
There is an unknown and persistent fundamental $\theta\in\mathbb{R}$. For each $t\in\mathcal{T}$, let $\theta_{t}$ be the state at time $t$ and for each $t>1$, [^13] For any $\alpha\in(0,1)$, analogs of our results apply for $\theta_{t}=\alpha\theta+(1-\alpha)\lambda_{t-1}$; see Online Appendix VI.
@@
Our law of motion for the endogenous state is simple. Nonetheless, we think it well-approximates a range of environments such as currency attacks or bank runs in which the failure probability depends on a single sufficient statistic: present reserves. This, in turn, is simply the sum of its initial reserves less the ‘current’ short/withdrawal positions. For a currency attack where the risky action ($a=1$) is to attack, initial reserves are $-\theta$ and the current short position is $\lambda_{t-1}$; for a bank run the risky action ($a=1$) is to deposit, initial reserves are $\theta$, and the current withdrawal position is $1-\lambda_{t-1}$. [^14] As is standard in global games, we have assumed that $\theta$ is supported on the real line, but it can be appropriately shifted in space and truncated so that the sign fits the economic phenomena in question.
@@
$u(a_{it},\theta_{t}):=\begin{cases}\theta_{t}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$
$\lambda_{0}$ is common knowledge but interim payoffs as well as $(\lambda_{t})_{t\geq 1}$ are unobserved and inferred via beliefs about the state $\theta$ [^15] An identical assumption is made in the context of dynamic global games by Angeletos, Hellwig, and Pavan (2007). Our model is equivalent to one in which agents observe noisy signals of past play $(\lambda_{t})_{t\geq 1}$, or both signals about the state as well as past play—we formalize this an equivalence in Appendix C and discuss this below.
Interpreting initial play. We interpret $\lambda_{0}\in[0,1]$ as a shock. This is equivalent to a one-off and transitory shock to the cost at time $1$. [^16] More explicitly, suppose there is zero initial play but there is an initial shock to costs of playing the risky action. Costs are potentially time-varying $c_{1},c_{2},\ldots$ with $c:=c_{2}=c_{3}=c_{4}=\ldots$ and $c_{1}\neq c$. Then, the dynamics of this model with zero initial play is equivalent to the model with $\lambda_{0}=c-c_{1}$. The impact of this shock is equivalent to that of non-zero initial play: positive shocks which incentive the risky action corresponds to high $\lambda_{0}$ and vice versa. In the language of macroeconomics, this corresponds to an unexpected shock (“MIT shock”) and we analyze the dynamics of aggregate play following the shock. In the language of fictitious play, this corresponds to a shock to initial play and we analyze how the speed of learning shapes the limit action profile.
Timing and Information. The timing of the game and the informational environment are as follows. All agents share a common improper uniform prior about $\theta$. At each time $t\in\mathcal{T}$, each agent $i$ receives a private signal $x_{it}\sim N(\theta,\sigma_{t}^{2})$ that is iid across agents and independent across time. Each agent $i$ then updates to her time-$t$ posterior about $\theta$ conditional on the natural filtration $\mathcal{F}_{t}$ generated by past signals $\{x_{is}\}_{s=1}^{t}$. From standard Bayesian updating, this posterior is also Gaussian: $\theta|\mathcal{F}_{t}\sim N(\mu_{it},\eta_{t}^{2})$, where $\mu_{it}:=\eta_{t}^{2}\sum_{s=1}^{t}\sigma_{s}^{-2}x_{is}$ and $\eta_{t}^{2}:=(\sum_{s=1}^{t}\sigma_{s}^{-2})^{-1}$. Finally, each agent $i$ chooses $a_{it}\in\{0,1\}$ to maximize her expected utility. Note that since agents are atomless, they cannot influence future aggregate play, so it is without loss to assume they are myopic. [^17] For example, we could redefine agent $i$’s payoffs as $\max_{(a_{it})_{t\geq 1}}\sum_{t\geq 1}\delta^{t}\mathbb{E}_{t}[u(a_{it},\lambda_{t-1},\theta)]$ for some $\delta\in(0,1)$.
@@
  • State transition & non-contemporaneous incentives: Our model is one in which the state at time-$t$ ($\theta_{t}$) is shaped by play at time-$t-1$ ($\lambda_{t-1}$). A few remarks are in order. First, we think this captures an important element of dynamic coordination phenomena. For instance, in currency crises the susceptibility of the current regime depends on the short position. Note also that our model exhibits path dependence—though $\theta_{t}$ depends directly only on $\lambda_{t-1}$, it also depends indirectly on the full path of past play $(\lambda_{s})_{s<t-1}$. [^20] This is broadly analogous to how an AR$(1)$ process depends indirectly on realizations of the process which are far into the past (although our setting is non-stationary). An alternate interpretation is that our model is a form of best-response dynamics. Of course, some care is required with this interpretation since past play is unobserved: players are thus best-responding to their best guess of past play. Experimental evidence in (complete information) coordination games find that the vast majority (96%) of play constitute myopic best-responses (Mäs and Nax, 2016). This allows our model to make rich and nuanced predictions on how the speed of learning shapes equilibrium selection. By contrast, a model with exact contemporaneous incentives cannot capture this—the risk-dominant equilibria is always selected. [^21] In Appendix B we show that the unique belief threshold is $c-1/2$ so the risk-dominant action is always selected in the limit. Second, our main result on learning regimes (Theorem 1) holds for vanishing inertia: if we fixed the speed of learning as a function of calendar time but shrink the spacing between time periods, the sub/super-quadratic threshold for risk/non-risk dominance continues to hold. This is in the same spirit as the ‘frictionless limit’ of Frankel and Pauzner (2000) where switching frictions vanish. Third, as the degree of inertia vanishes, the (unique) equilibrium of our inertial coordination game converges to an equilibrium of the coordination game with contemporaneous incentives. [^22] That is, with utility function $u(1,\theta,\lambda_{t})=\theta+\lambda_{t}$. Proposition 4 in Appendix B formalizes this. This means that the speed of learning relative to the degree of inertia is what is important for obtaining richer predictions about equilibrium outcomes.
@@
  • Observability. We have assumed agents do not observe aggregate past play. [^23] Aggregate past play is an injective function of the state, so if agents were to observe this, they would immediately learn the state. However, our model is equivalent to one in which agents receive private signals both about the state, and about aggregate past play. [^24] For instance, see Dasgupta (2007); Trevino (2020). Explicitly, suppose that agent $i\in[0,1]$, in addition to observing $x_{it}\sim N(\theta,\sigma^{2}_{t})$, additionally observes an independent signal distributed $y_{it}\sim N(\Phi^{-1}(\lambda_{t-1}),\tau_{t}^{2})$; this signal structure was introduced by Dasgupta (2007) and used recently by Trevino (2020) to study financial contagion. Then, we can show by induction that equilibrium behavior is equivalent to a model in which players receive a more precise signal about $\theta$ and no information about aggregate past play; Appendix C formalizes this connection. In light of this observation, we load all learning on the state, and agents make inferences about aggregate past play conditional on their beliefs about the state.
@@
  • Continuum of players. We follow the literature on global games by assuming that there is a continuum of atomless players. By the continuum law of large numbers, this implies that conditioned on the fundamental state $\theta$, all aggregate randomness in private signal realizations across the population washes out. This allows us to analyze the dynamics of aggregate play cleanly as the solution to a non-linear deterministic difference equation. In Online Appendix IV we develop a finite-player version of our model with $N<+\infty$ players and show via standard concentration arguments of the empirical distribution of beliefs that limit play behaves similarly in the finite-player version of our model (see Theorem 2 in Online Appendix IV).

### v3

@@ 1 Introduction
Our main result (Theorem 1) develops a tight connection between the speed of learning and limit play: the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically. Sub-quadratic learning includes canonical learning environments such as (i) independent and identically distributed signals; and (ii) the receipt of a one-time signal. On the other hand, super-quadratic learning includes social learning such that, in each period, agents are randomly matched and sample their partner’s information sets. The tight connection we develop between learning speeds and limit play has simple but sharp economic implications. Slow (sub-quadratic) learning implies history independence: the initial shock to aggregate play as well the fine details of what agents learn in the interim is irrelevant for limit play. Conversely, when learning speeds are fast (super-quadratic), this induces history dependence such that initial shocks can be self-fulfilling: whether shocks propagate or not is jointly sensitive to both the fundamentals and the magnitude of the shock. [^2] See Section I of Krugman (1991) for a discussion of history-dependence and expectations in economics.
@@
Our model of inertial coordination games imposes a lag between players’ actions and their effect on the time-varying state. Hence, when a player contemplates whether or not to play the risky action, his belief about the current state depends on his belief about past play. But that, in turn, depends on his belief about other players’ past beliefs which themselves are about play yet further back into the past. Thus, the path of past play (and beliefs)—even of the distant past—multiply forward to shape incentives in the present. [^3] Note that our environment is, in a loose sense, the opposite of coordination games with switching frictions (Frankel and Pauzner, 2000). In those models, switching frictions lead players to form beliefs about future play and hence future beliefs; by contrast, in our model, inertia leads players to form beliefs about past play and hence past beliefs. An important by-product of inertia is that it offers a simple yet novel means of obtaining unique predictions in coordination environments.
At the same time, our model is motivated by—and consistent with—recent work documenting empirical facts about bank runs. In their empirical analysis of minute-by-minute depositor-level data from an Indian bank, Iyer and Puri (2012) find that depositors’ beliefs about “the fraction of other depositors… that have run until time t1t-1” help determine whether they run at time $t$ (eg, inertia). More generally, Correia, Luck, and Verner (2024) use historical data from American commercial banks to assess the two prevailing explanations of bank failure in this literature: the “bank runs view” in which failures are due to self-fulfilling runs on solvent banks, and the “solvency view” in which insolvency due to poor fundamentals is the primary reason underlying bank failure. In addition to matching Correia, Luck, and Verner (2024)’s finding that “depositors tend to be slow to react to an increased risk of bank failure” (eg, the inertia), our main result suggests that which of the above two explanations is best is characterized by the speed of learning among depositors: slow learning implies that outcomes are solely determined by fundamentals as in the “solvency view,” while fast learning implies that self-fulfilling runs may occur even for banks with strong fundamentals as in the “bank runs view.” Hence, one of our contributions is to offer a simple dynamic model which reconciles empirical regularities about ‘real world’ coordination environments with the more abstract logic of learning and higher-order beliefs in coordination games. Further, our characterization of the connection between speed and limit play is invariant to the degree of inertia (duration of lag) and, as inertia vanishes, the equilibrium converges to an equilibrium of the model with purely contemporaneous incentives. [^4] These claims are made precise in Appendix II.
@@
In what sense does our model relate to the extant literature on coordination games? It turns out that static global games (Carlsson and Van Damme, 1993) are recovered exactly as a special case of our model in which players only receive a single signal at time $t=1$: since this learning process is sub-quadratic (posterior precisions are unchanged after $t=1$), the risk-dominant profile is—as in global games—played in the limit. This connection arises from observing that when there is no interim learning, best-responding to expected past play along each time step in our model is equivalent to the interim elimination of strictly dominated strategies in the static game. As we make precise in Appendix V, this uncovers a formal connection between “contagion” arguments over type spaces (which encode hierarchies of higher-order beliefs) and “contagion” over time (where the current state depends on past play). Thus, our results generalize the predictions of static global games to sub-quadratic learning and offer a simple dynamic learning foundation for the ‘vanishing noise limit’ commonly invoked in that literature.
@@
Finally, our model makes sharp qualitative predictions about the path of play (Proposition 3) whenever learning is slow enough such that the risk-dominant profile is played in the limit. In particular, transition dynamics toward limit play depend starkly on the precision of signals. When signals are precise, aggregate play exhibits a sudden phase transition from approximately all players playing the non-risk-dominant action (potentially for a long while) to approximately all players playing the risk-dominant action thereafter. Conversely, when signals are noisy, aggregate play exhibits a gradual transition as the measure of players playing risk-dominant action increases gradually. These differing transition regimes are driven by differences in time-varying heterogeneity in beliefs. The importance of heterogeneity has been recognized as a crucial determinant of equilibrium uniqueness in coordination games (Morris and Shin, 2006). Our results complement this insight by highlighting that the speed of learning—which drives how the cross-sectional distribution of beliefs vary over time—is also a crucial determinant of the path of play. It also suggests that “spikes” in the time-series of aggregate actions (e.g., withdrawal decisions or adoption of networked goods) can be consistent with transition to limit equilibrium play and need not necessarily correspond to “equilibrium shifts” (Chwe, 2013; Morris and Yildiz, 2019).
@@
Related literature. Our paper directly relates to the large literature on global games. A broad takeaway from the literature on static global games (Carlsson and Van Damme, 1993; Kajii and Morris, 1997; Frankel, Morris, and Pauzner, 2003, among others) is that perturbing common knowledge by introducing a vanishing amount of private uncertainty often selects the risk dominant equilibrium. This approach has been widely deployed in a variety of economic applications to select among multiple equilibria (see, e.g., Morris and Shin (1998); Goldstein and Pauzner (2005) among many others). Our model nests the canonical global games model of Carlsson and Van Damme (1993) as a special case with a private signal at $t=1$ and no learning thereafter. The “contagion argument” in global games where players iteratively eliminate interim dominated strategies is sometimes viewed metaphorically or as a convenient way to select among equilibria. [^5] See, Morris (2002) and Chapter 2 of Spiegler (2024). Our simple dynamic environment with an endogeneously changing state thus offers an explicit dynamic foundation for the “contagion” logic of global games consistent with recently documented facts on bank runs (Iyer and Puri, 2012; Correia, Luck, and Verner, 2024). Moreover, we substantially generalize the prediction of risk-dominant selection to the regime of sub-quadratic learning.
@@ 2 Model
Payoffs. There is a unit measure of agents indexed by $i\in[0,1]$. [^12] Endow $[0,1]$ with the Lebesgue extension constructed in Sun and Zhang (2009) so that the continuum law of large numbers holds. Time is discrete and indexed by $t\in\mathcal{T}=\{1,2,\ldots\}$. At each time $t\in\mathcal{T}$, each agent $i$ chooses between a risky action $(a_{it}=1)$ or a safe action $(a_{it}=0)$. Let $\lambda_{t}$ denote the measure of agents who play the risky action at time $t$.
There is an unknown and persistent fundamental $\theta\in\mathbb{R}$. For each $t\in\mathcal{T}$, let $\theta_{t}$ be the state at time $t$ and for each t>1t>1, [^13] For any $\alpha\in(0,1)$, analogs of our results apply for $\theta_{t}=\alpha\theta+(1-\alpha)\lambda_{t-1}$; see Appendix VI.
@@
Our law of motion for the endogenous state is simple. Nonetheless, we think it well-approximates a range of environments such as currency attacks or bank runs in which the failure probability depends on a single sufficient statistic: present reserves. This, in turn, is simply the sum of its initial reserves less the ‘current’ short/withdrawal positions. For a currency attack where the risky action ($a=1$) is to attack, initial reserves are $\theta$ and the current short position is $\lambda_{t-1}$; for a bank run the risky action ($a=1$) is to deposit, initial reserves are $\theta$, and the current withdrawal position is $1-\lambda_{t-1}$. [^14] As is standard in global games, we have assumed that $\theta$ is supported on the real line, but it can be appropriately shifted in space and truncated so that the sign fits the economic phenomena in question.
@@
[…]
In the main text, we assume that $\lambda_{0}$ is common knowledge but interim payoffs as well as $(\lambda_{t})_{t\geq 1}$ are unobserved and inferred via beliefs about the state $\theta$. [^15] An identical assumption is made in the context of dynamic global games by Angeletos, Hellwig, and Pavan (2007). As Appendix VI.1 shows, the assumption that $\lambda_{0}$ is common knowledge may be relaxed. Our model is equivalent to one in which agents observe noisy signals of past play $(\lambda_{t})_{t\geq 1}$, or both signals about the state as well as past play—we formalize this an equivalence in Appendix III and discuss this below.
Interpreting initial play. We interpret $\lambda_{0}\in[0,1]$ as a shock. This is equivalent to a one-off and transitory shock to the cost at time 11.161616More explicitly, suppose there is zero initial play but there is an initial shock to costs of playing the risky action. Costs are potentially time-varying $c_{1},c_{2},\ldots$ with $c:=c_{2}=c_{3}=c_{4}=\ldots$ and $c_{1}\neq c$. Then, the dynamics of this model with zero initial play is equivalent to the model with $\lambda_{0}=c-c_{1}$. The impact of this shock is equivalent to that of non-zero initial play: positive shocks which incentivize the risky action correspond to high $\lambda_{0}$ and vice versa. In the language of macroeconomics, this corresponds to an unexpected shock (“MIT shock”) and we analyze the dynamics of aggregate play following the shock. In the language of fictitious play, this corresponds to a shock to initial play and we analyze how the speed of learning shapes the limit action profile.
Timing and Information. The timing of the game and the informational environment are as follows. All agents share a common improper uniform prior about $\theta$. At each time $t\in\mathcal{T}$, each agent $i$ receives a private signal $x_{it}\sim N(\theta,\sigma_{t}^{2})$ that is iid across agents and independent across time. Each agent $i$ then updates to her time-$t$ posterior about $\theta$ conditional on the natural filtration $\mathcal{F}_{it}$ generated by past signals $\{x_{is}\}_{s=1}^{t}$. From standard Bayesian updating, this posterior is also Gaussian: $\theta|\mathcal{F}_{it}\sim N(\mu_{it},\eta_{t}^{2})$, where $\mu_{it}:=\eta_{t}^{2}\sum_{s=1}^{t}\sigma_{s}^{-2}x_{is}$ and $\eta_{t}^{2}:=(\sum_{s=1}^{t}\sigma_{s}^{-2})^{-1}$. Finally, each agent $i$ chooses $a_{it}\in\{0,1\}$ to maximize her expected utility. Note that since agents are atomless, they cannot influence future aggregate play, so it is without loss to assume they are myopic. [^17] For example, we could redefine agent $i$’s payoffs as $\max_{(a_{it})_{t\geq 1}}\sum_{t\geq 1}\delta^{t}\mathbb{E}_{t}[u(a_{it},\lambda_{t-1},\theta)]$ for some $\delta\in(0,1)$.
@@
  • State transition & non-contemporaneous incentives: Our model is one in which the state at time-$t$ ($\theta_{t}$) is shaped by play at time-t1t-1 ($\lambda_{t-1}$). A few remarks are in order. First, we think this captures an important element of dynamic coordination phenomena (such as the bank runs studied in Iyer and Puri (2012) and Correia, Luck, and Verner (2024)). For instance, in currency crises the susceptibility of the current regime depends on the short position. Note also that our model exhibits path dependence—though $\theta_{t}$ depends directly only on $\lambda_{t-1}$, it also depends indirectly on the full path of past play $(\lambda_{s})_{s<t-1}$. [^20] This is broadly analogous to how an AR(1)(1) process depends indirectly on realizations of the process which are far into the past (although our setting is non-stationary). An alternate interpretation is that our model is a form of best-response dynamics. Of course, some care is required with this interpretation since past play is unobserved: players are thus best-responding to their best guess of past play. Experimental evidence in (complete information) coordination games find that the vast majority (96%) of play constitute myopic best-responses (Mäs and Nax, 2016). This allows our model to make rich and nuanced predictions on how the speed of learning shapes equilibrium selection. By contrast, a model with exact contemporaneous incentives cannot capture this—the risk-dominant equilibrium is always selected. [^21] In Appendix II we show that the unique belief threshold is c1/2c-1/2 so the risk-dominant action is always selected in the limit. Second, our main result on learning regimes (Theorem 1) holds for vanishing inertia: if we fixed the speed of learning as a function of calendar time but shrink the spacing between time periods, the sub/super-quadratic threshold for risk/non-risk dominance continues to hold. This is in the same spirit as the ‘frictionless limit’ of Frankel and Pauzner (2000) where switching frictions vanish. Third, as the degree of inertia vanishes, the (unique) equilibrium of our inertial coordination game converges to an equilibrium of the coordination game with contemporaneous incentives. [^22] That is, with utility function $u(1,\theta,\lambda_{t})=\theta+\lambda_{t}$. Proposition 4 in Appendix II formalizes this. This means that the speed of learning relative to the degree of inertia is what is important for obtaining richer predictions about equilibrium outcomes.
@@
  • Observability. We have assumed agents do not observe aggregate past play. [^23] Aggregate past play is an injective function of the state, so if agents were to observe this, they would immediately learn the state. However, our model is equivalent to one in which agents receive private signals both about the state, and about aggregate past play. [^24] For instance, see Dasgupta (2007); Trevino (2020). Explicitly, suppose that agent $i\in[0,1]$, in addition to observing $x_{it}\sim N(\theta,\sigma^{2}_{t})$, additionally observes an independent signal distributed $y_{it}\sim N(\Phi^{-1}(\lambda_{t-1}),\tau_{t}^{2})$; this signal structure was introduced by Dasgupta (2007) and used recently by Trevino (2020) to study financial contagion. Then, we can show by induction that equilibrium behavior is equivalent to a model in which players receive a more precise signal about $\theta$ and no information about aggregate past play; Appendix III formalizes this connection. In light of this observation, we load all learning on the state, and agents make inferences about aggregate past play conditional on their beliefs about the state.
@@
  • Continuum of players. We follow the literature on global games by assuming that there is a continuum of atomless players. By the continuum law of large numbers, this implies that conditioned on the fundamental state $\theta$, all aggregate randomness in private signal realizations across the population washes out. This allows us to analyze the dynamics of aggregate play cleanly as the solution to a non-linear deterministic difference equation. In Appendix IV, we develop a finite-player version of our model with $N<+\infty$ players and show via standard concentration arguments of the empirical distribution of beliefs that limit play behaves similarly in the finite-player version of our model (see Theorem 2 in Appendix IV).